A consumer spend all his income ( in equilibrium) to purchase 100 units of X and 200 units of Y when price of X and Y is Rs.2 and Rs.4 per unit respectively. If the price of Y is decrease to Rs.4 and the consumer increase his consumption of X to 200 units determine the own price and cross price elasticity of demand for X
1
Expert's answer
2020-08-26T12:59:26-0400

If the price of Y decreases from Rs.4 to Rs.2 and the consumer increases his consumption of X to 200 units, then we can't determine the own price elasticity, because we don't know the change in price for X.

Cross price elasticity of demand for X is:

"Ed = \frac{200 - 100} {4 - 2} × \frac{4 + 2} {200 + 100} = 1."

So, the goods X and Y are substitutes.